# To Offload or Not To Offload: Model-driven Comparison of Edge-native and On-device Processing

### v2

@@ III-E Discussion
Generalizing to other accelerator-driven workloads. While the above analysis focuses on DNN workloads with deterministic service times modeled using an $M/D/1$ system, our modeling framework can be extended to other workloads with different service time characteristics by incorporating appropriate queuing results. For example, workloads with recurrent architectures such as GRUs and LSTMs often process variable-length inputs with service times that vary with the input sequence length. Similarly, early-exit DNNs allow requests to exit at different layers based on confidence thresholds, which results in variable service times. To capture such variability, our framework can adopt an $M/M/1$ queuing system where service times follow an exponential distribution with the expected wait time shown in (6). By selecting the appropriate queuing model, our framework can adapt to diverse workloads and provide accurate response time estimates as will be shown in Section IV-D. Currently, our framework does not account for workloads that violate the FCFS assumption, such as Large Language Model (LLM) inference with continuous batching where requests are dynamically batched, and extending the models to capture such behavior is a subject of future work.
@@
Model limitations. Our analysis assumes request arrivals follow a Poisson process and are independent, which may not hold under bursty arrivals. To address this, our models can adopt a $G/G/1$ system, which accommodates arbitrary arrival and service time distributions. Although no closed-form solution exists for the expected wait time in this setting, Marshall et al. [40] provide an upper bound {dmath} E[w_G/G/1] ≤λ(σa2+ σs2)2(1-ρ) where $\sigma_{a}$ and $\sigma_{s}$ are the variances for the interarrival and service distributions respectively. Additionally, the queuing results used in our analysis assume that each hardware queue has infinite capacity. For scenarios with bounded queue sizes, we can incorporate results for finite-buffer queuing systems [19] to model the impact of such capacity constraints.
IV Model Validation
In this section, we empirically evaluate how the key factors modeled in our analytic framework affect the performance of the two execution approaches and demonstrate that our models can accurately predict the response times across a wide range of configurations.
IV-A Experimental Setup
@@
Workloads. We validate our model’s ability to capture the impact of workload characteristics using a diverse set of workloads, including DNN-based models as well as recurrent models. For DNN workloads, we focus on cognitive assistant applications involving image classification and object detection. Specifically, we use three widely deployed DNN architectures, namely MobileNet, Inception, and YOLO, each with distinct computational demands and input sizes as shown below:
@@
All inference requests are executed using TensorFlow 2.13.0 across the evaluated hardware. We use frames extracted from a 720p dash cam video as input to the models, with each frame processed either locally on the device or offloaded to the edge. Since application workloads are often dynamic (e.g., triggered by motion detection), we implement a workload generator that selects video frames and generates requests following a Poisson process, which runs on a separate machine to avoid performance interference with the device during evaluation. For recurrent models, we focus on personal voice assistant applications involving sentiment classification of conversational utterances. We validate our models using two types of recurrent neural networks: a GRU model with lower computational demand and a more compute-intensive LSTM model, with their architectural specifications summarized below:
Model Embedding Dim # Layers $\times$ Units Activation
GRU 512 2 $\times$ 512 Sigmoid
LSTM 1024 3 $\times$ 1024 Sigmoid
Both models are implemented using TensorFlow’s Keras API. To simulate real-world user inputs, we use the Amazon Alexa Topical Chat dataset [41], a knowledge-grounded human-to-human conversation dataset consisting of spoken-style dialogues with an average utterance length of 19 tokens.
@@
Hardware. We use a diverse set of hardware platforms to reflect real-world deployment scenarios. For client devices with accelerators, we use the NVIDIA Jetson TX2 and Jetson Orin Nano, two commonly deployed platforms in IoT applications, with the TX2 being an earlier model and the Orin Nano offering better performance. For edge servers, we use a Dell PowerEdge R630 equipped with an NVIDIA A2 GPU, a 40-core Intel Xeon E5-2660v3 CPU, and 256GB RAM. We also use a server with an NVIDIA RTX 4070 GPU, 8-Core AMD Ryzen 7 5700X CPU, and 32GB RAM. Both edge servers are connected to a 1 Gbps network. These platforms span a wide performance range, from 1.3 TFLOPS (TX2) to 2.6 TFLOPS (Orin Nano), 4.5 TFLOPS (A2), and up to 29.1 TFLOPS (RTX 4070) in peak FP16 throughput. For each accelerator, the effective degree of parallelism $k$ depends on both the number of cores and the workload’s processing and memory demands, and is not directly observable. To estimate $k$, we empirically measure how response time varies with request rate for each workload and identify a value of $k$ that best captures the observed scaling behavior. All platforms run Ubuntu 18.04, and we use the Linux Traffic Control (TC) subsystem to emulate different network bandwidths.
IV-B Parameter Estimation
@@
Our model requires several parameters to compute the end-to-end response times. We now describe how these parameters are obtained in our evaluation and outline other practical methods for estimating them.
@@
Estimating average service times. Our system estimates workload service times through empirical profiling and a neural network–based prediction approach. We estimate service times for DNN workloads by monitoring inference durations using the NVIDIA System Management Interface (nvidia-smi) tool [23], which reports per-process execution times on the accelerator. For GRU and LSTM workloads, we profile request latencies on a representative input set distinct from the test set and use the average latency from this profiling as the service time of the model. This profiling data is used for our edge-offloading and on-device processing experiments. For collaborative processing, a neural network is trained using profiling data to predict service time based on prior work discussed in Section III-B. Such a neural network approach avoids the extensive profiling overhead of all possible split configurations by predicting partial execution service times.
@@
Estimating average network delays. Estimating transmission delay requires knowledge of the available network bandwidth, which we continuously measure in real-time using iperf [42]. In addition to queuing and transmission delays captured by our models, network delays also include propagation and processing delays. Since we assume that requests are offloaded to a nearby edge server, the propagation delay (caused by signals traveling at the speed of light) is minimal and thus excluded from our models. Processing delay can be estimated by measuring the round-trip time of small probe packets, and is excluded from our models due to its negligible impact on overall latency.
@@
IV-C Impact of Workload Characteristics
@@
Our first experiment investigates how workload characteristics influence offloading decisions using three DNN workloads with distinct profiles: 1) MobileNetV2 (low compute, small payload); 2) InceptionV4 (high compute, moderate payload); and 3) YOLOv8n (high compute, large payload). We compare on-device processing on TX2 and Orin Nano against offloading to an A2 GPU using a request rate of 2 requests per second (RPS) and a network speed of 5 Mbps. For each device, we test all available GPU frequency settings to emulate different device conditions.
Figure 3: Latency comparison across workloads at 5 Mbps network speed and 2 RPS. (a-b) MobileNetV2, (c-d) InceptionV4, and (e-f) YOLOv8n. Left: TX2 vs A2; Right: Orin Nano vs A2.
@@
Figure 3 compares the average latency of on-device processing and edge offloading across workloads, devices, and GPU frequency levels. Figures 3a and 3b show that for MobileNetV2, local processing on both devices outperform offloading: TX2 achieves lower latency in its three highest frequency modes, while Orin Nano consistently outperforms offloading across all frequency levels. The trend reverses for InceptionV4 as shown in Figures 3c and 3d, where offloading provides lower average latency across all device configurations. Lastly, Figures 3e and 3f shows the results for YOLOv8n. In contrast to the results for InceptionV4, TX2 outperforms offloading in its top six performance modes despite YOLOv8n’s high compute demand, while Orin Nano achieves lower average latency than offloading across all performance modes. We observe similar results when offloading to RTX4070 and omit the details here due to space constraints. In all cases, the model predictions closely match the observed response times for both approaches, achieving a mean absolute percentage error of 2.2%, with 91.5% of predictions falling within $\pm$5of the observed latency and 100% within $\pm$10%, proving the accuracy of our models.
@@
Collectively, these empirical results align with the qualitative observations stated in Remark III.1 and Remark III.2. While the performance ratio between edge and device is fixed, the absolute advantage of offloading scales with the workload’s demands. For workloads with low computational demand (MobileNetV2), the edge’s relative speed advantage results in only small absolute gains that can be overshadowed by network delays of offloading. However, for compute-intensive workloads (InceptionV4), the same performance ratio yields substantially larger absolute improvements that outweigh the network delays. Moreover, workloads with large message payloads (YOLOv8n) favor on-device processing because it does not involve any network components and thus avoids any network delay.
@@
Key takeaway. Offloading provides lower latency when the edge’s absolute performance gain exceeds the associated network delay. Our models accurately predict the latency for both strategies with a mean absolute percentage error of 2.2%.
IV-D Impact of Complex Deep Learning Models
@@
Figure 4: Latency comparison across execution strategies for (a) GRU model and (b) LSTM model at 3 RPS under 10 Mbps.
Figure 5: Latency comparison across execution strategies for MobileNetV2 at 2 RPS with varying network bandwidth: (a) RTX4070 as the edge server, (b) A2 as the edge server.
Figure 6: Latency comparisons across (a) different collaborative processing configurations, (b) request rates under 10 Mbps and 20 Mbps networks, and (c) numbers of co-located applications.
@@
In addition to the DNN workloads, our model also accurately captures the latency behavior of more complex deep learning workloads. Figure 4 compares the average latency of on-device processing on the Orin Nano against offloading to A2 for two recurrent models: GRU and LSTM, under a moderate load of 3 RPS. Since service times vary with input length in these models, we incorporate the $M/M/1$ queuing model into our analysis to account for this variability in latency prediction.
Figure 4a shows that even for the lightweight GRU model, offloading offers lower latency due to the small input utterance length (on average less than 100 bytes), which incurs minimal network delay. Figure 4b shows that as the model complexity increases, the absolute performance gain of offloading increases, leading to a greater latency reduction. In both cases, our model predictions closely match the observed latencies.
@@
Key takeaway. Our models generalize to complex deep learning workloads beyond DNN inference by integrating appropriate queuing models.
IV-E Impact of Network Conditions
Figure 5 shows how network speed affects the effectiveness of each execution strategy. Figure 5a compares the latency of on-device processing on TX2 and Orin Nano with edge offloading to RTX4070 under varying network speeds. For both devices, we observe a performance crossover point where offloading becomes faster than on-device processing. The location of the crossover depends on both network bandwidth and the performance gap between the device and the edge. For the lower-performance TX2, offloading becomes advantageous at 5 Mbps, while the faster Orin Nano requires a higher bandwidth (15 Mbps) for offloading to be beneficial.
@@ IV-E Impact of Network Conditions
Figure 5b shows that even for a single device, multiple crossover points may exist depending on network conditions and the device’s performance. For example, at 5 Mbps, TX2 processing in its top three performance modes outperforms offloading; however, as TX2’s performance decreases beyond those modes, offloading becomes more advantageous. Across all scenarios, our models closely predict the crossover points, demonstrating their accuracy in capturing the tradeoff between the computation speedup of edge offloading and network delay.
@@
Key takeaway. At high network bandwidths, edge offloading becomes favorable as its computational advantage outweighs the reduced network delays.
IV-F Impact of Collaborative Processing
@@
To demonstrate our models’ ability to capture collaborative (i.e., split) processing behavior, we consider an example using MobilNetV2 and use the Orin Nano as the device and A2 as the edge under a 50 Mbps network. We evaluate three split points (SPs) where the model is progressively offloaded to the edge and compare them to full on-device execution. For each split point, the initial portion of the neural network is executed locally on the device, while the resulting intermediate output and remaining computation are offloaded to the edge. Figure 6a plots the observed latencies of each strategy alongside model predictions. To compute the end-to-end latency, we use the combined analytic models for device and edge processing as discussed in Section III-E. As shown, the latency increases with the amount of compute offloaded. This is because the intermediate results of later layers grow in size, adding to the network transfer overhead. As seen, for MobileNetV2 and the chosen network configuration, local processing is faster than split processing. Across all split points, the model-predicted latencies closely align with observed values. Note that other models will exhibit other types of behaviors with split processing, which we omit due to space constraints.
@@
Key takeaway. Our models can be extended to capture both compute and network delays in device-edge collaborative processing.
IV-G Impact of Request Rate
@@
Figure 6b shows how the request rate $\lambda$ affects offloading decisions using the MobileNetV2 workload under 10 Mbps and 20 Mbps network bandwidth, with both the TX2 and Orin Nano set to their maximum performance modes. We observe two distinct behaviors depending on network bandwidth. At 10 Mbps, on-device processing using Orin Nano outperforms offloading across all RPS. However, at 20 Mbps, a performance crossover occurs: on-device processing is faster at low load (RPS $\leq$ 60), while offloading becomes more efficient beyond that point. In contrast, processing on TX2 consistently exhibits higher latency across all RPS due to its limited performance. In all cases, as the system approaches saturation at high RPS, queue lengths show increased variability from the Poisson arrival process, leading to greater fluctuations in queuing delays and response times. Overall, the predicted end-to-end latencies of our models match closely with the observed latency within confidence intervals, demonstrating the accuracy of our models.
This result validates the competing scaling effects of $\lambda$ discussed in Lemma III.1. Although both the network queuing delays (for offloading) and local processing queuing delay (for on-device processing) grow with $\lambda$, the rate at which the delays increase differs depending on the hardware and network configurations. At 10 Mbps, the network queuing delay grows faster than the device queuing delay as $\lambda$ increases, favoring on-device processing. Conversely, at 20 Mbps, device queuing delays become dominant at higher $\lambda$, shifting the advantage toward offloading.
@@
Key takeaway. As load increases, on-device processing becomes more advantageous under slower networks, while offloading provides lower latency on faster networks.
IV-H Impact of Multi-Tenancy
We now experimentally validate the accuracy of our models in multi-tenant settings, using Orin Nano as the device and A2 as the edge server. We vary the number of applications deployed on the edge, each serving InceptionV4 requests, with all models loaded in memory concurrently. Each application receives 2 RPS from its corresponding client. We report the average offloading latency and compare it with on-device processing.

### v3

@@ 3.5. Discussion
Generalizing to other accelerator-driven workloads. While the above analysis focuses on DNN workloads modeled with deterministic service times, our models can be extended to other accelerator-driven workloads with different service time characteristics by incorporating appropriate queuing formulations. For example, recurrent models have variable execution times that depend on input length. Similarly, LLM inference exhibits variable service times due to variations in the number of autoregressive decoding steps. To capture such variability, our models can adopt an $M/M/1$ system in which service times follow an exponential distribution, with the expected wait time given in (7). By selecting the appropriate queuing model, our models can adapt to diverse workloads and provide accurate response time estimates as shown in Sec 4.4.
@@
Model limitations. Our analysis assumes request arrivals follow a Poisson process and are independent, which may not hold under bursty arrivals. To address this, our models can adopt a $G/G/1$ system, which accommodates arbitrary arrival and service time distributions. Although no closed-form solution exists for the expected wait time in this setting, Marshall et al. (marshall1968) provide an upper bound
(12) $E[w_{G/G/1}]\leq\frac{\lambda(\sigma_{a}^{2}+\sigma_{s}^{2})}{2(1-\rho)}$
where $\sigma_{a}^{2}$ and $\sigma_{s}^{2}$ are the variances of the request interarrival time and service time distributions, respectively. Additionally, the queuing results used in our analysis assume that each queue has infinite capacity. For scenarios with bounded queue sizes, we can incorporate results for finite-buffer queuing systems (Harchol2013) to model the impact of such capacity constraints.
4. Model Validation
In this section, we empirically validate the accuracy of our analytic models for predicting the end-to-end response times of edge offloading, on-device processing, and collaborative processing across a wide range of configurations.
4.1. Experimental Setup
@@
Workloads. We validate our model’s ability to capture the impact of workload characteristics using DNN, RNN, and LLM workloads. For DNN workloads, we focus on inference tasks for AR applications, specifically image classification, using three widely deployed architectures: MobileNet, Inception, and YOLO, each with distinct compute demands and input sizes:
@@
All requests are executed using TensorFlow 2.13.0 on frames extracted from a 720p video. For RNN, we use a Long Short-Term Memory (LSTM) model implemented using TensorFlow’s Keras API for sentiment classification, with the Amazon Alexa Topical Chat dataset (Topical2023) as input. For LLM, we use the Llama-3.2-1B model (llama3_2_1B) run via llama.cpp (llama_cpp), with prompts from the Hugging Face ShareGPT dataset (sharegpt2024) as input. The specifications of both models are summarized below:
Model Architecture Parameters # of Layers
LSTM Recurrent 40 M 3
Llama-3.2-1B Transformer 1.2 B 16
Since application workloads are often dynamic (e.g., triggered by motion detection), we implement a workload generator that generates requests following a Poisson process, running on a separate machine to avoid performance interference with the device.
@@
Hardware. For client devices, we use the NVIDIA Jetson TX2 and Jetson Orin Nano, two commonly deployed platforms in IoT applications, with the former being an earlier model and the latter offering higher performance. For edge servers, we use a Dell PowerEdge R630 equipped with an NVIDIA A2 GPU and 256GB RAM. We also use a server with an NVIDIA RTX 4070 GPU and 32GB RAM. Both edge servers are connected to a 1 Gbps network. These platforms span a wide performance range, from 1.3 TFLOPS (TX2) to 2.6 TFLOPS (Orin Nano), 4.5 TFLOPS (A2), and up to 29.1 TFLOPS (RTX 4070) in peak FP16 throughput. All platforms run Ubuntu 18.04. We use the Linux Traffic Control (TC) subsystem to emulate different network bandwidths. For each accelerator, the effective degree of parallelism $k$ depends on both the number of cores and the workload’s processing and memory demands, and is not directly observable. We estimate $k$ by empirically measuring how response time varies with request rate for each workload and identify a value of $k$ that best captures the observed scaling behavior.
4.2. Parameter Estimation
@@
Our models require several parameters to compute the end-to-end response times. We now describe how they are obtained in our evaluation. Additional practical approaches for estimating them are discussed in Sec 3.2.
@@
Estimating average service times. We estimate service times through profiling and a neural network–based prediction approach. For DNN workloads, we monitor inference durations using the NVIDIA System Management Interface (nvidia-smi) tool (nvidiasmi), which reports per-process execution times on the accelerator. For RNN and LLM workloads, we profile request latencies on a representative input set separate from the test set and use the average latency as input to the models. This profiling data informs our edge-offloading and on-device processing experiments. For collaborative processing, a neural network is trained on profiling data to predict service time as discussed in Sec 3.2 to avoid the extensive profiling overhead of all possible split configurations by predicting partial service times.
@@
Estimating average network delays. We estimate transmission delay using the available network bandwidth, measured continuously in real time with iperf (iperfwebsite). In addition to queuing and transmission delays captured by our models, network delays also include propagation and processing delays. Since we assume that requests are offloaded to a nearby edge server, the propagation delay (caused by signals traveling at the speed of light) is minimal and thus excluded from our models. Processing delay can be estimated by measuring the round-trip time of small probe packets, and is excluded from our models due to its negligible impact on overall latency.
@@
4.3. Impact of Workload Characteristics
@@
We first study how workload characteristics influence offloading decisions using three DNN models: 1) MobileNetV2 (low compute, small payload); 2) InceptionV4 (high compute, moderate payload); and 3) YOLOv8n (high compute, large payload). We compare on-device processing on TX2 and Orin Nano against offloading to an A2 GPU using a request rate of 2 requests per second (RPS) and a network speed of 5 Mbps.
Figure 2. Latency comparison for DNN workloads: (a-b) MobileNetV2, (c-d) InceptionV4, and (e-f) YOLOv8n.
@@
Figure 2 compares the average latency of on-device processing and edge offloading across workloads and devices. For each device, we test all available GPU frequency settings to emulate varying device conditions, with each data point representing one setting. Figures 2a and 2b show that for MobileNetV2, TX2 achieves lower average latency in its top three frequency levels, while Orin Nano consistently outperforms offloading across all frequency levels. The trend reverses for InceptionV4 as shown in Figures 2c and 2d, where offloading provides lower average latency across all device configurations. Lastly, Figures 2e and 2f show that despite YOLOv8n’s high compute demand, TX2 outperforms offloading in its top six frequency levels, while Orin Nano achieves lower average latency across all settings. We observe similar results when offloading to an RTX4070 GPU and omit the details here due to space constraints.
@@
Collectively, these results align with the qualitative observations stated in Remark 3.2 and Remark 3.3. While the performance ratio between device and edge is fixed, the absolute advantage of offloading scales with the workload’s computational demands. For MobileNetV2 (low compute), the edge’s relative speed advantage results in only small absolute gains that can be overshadowed by network delays of offloading. However, for InceptionV4 (high compute), the same performance ratio yields larger absolute improvements that outweigh the network delays. Moreover, YOLOv8n (large payloads) favors on-device processing because it avoids any network delay. In all cases, the model predictions closely match the observed response times for both approaches, achieving a mean absolute percentage error of 2.2%, with 91.5% of predictions falling within $\pm$5% of the observed latency and 100% within $\pm$10%, proving their accuracy.
@@
Key takeaway. Offloading provides lower average latency when the edge’s absolute performance gain exceeds the associated network delay. Our models accurately predict the latency for both strategies with a mean absolute percentage error of 2.2%.
4.4. Impact of Complex Deep Learning Models
@@
Figure 3. Latency comparison of execution strategies for (a) LSTM and (b) Llama-3.2-1B models.
@@
In addition to DNN workloads, our model accurately captures the latency behavior of other deep learning workloads. Figure 3 compares the average latency of on-device processing on the Orin Nano against offloading to the A2 GPU for RNN and LLM workloads under moderate loads. Since service times vary across requests in these models, we use the $M/M/1$ queuing model to account for this variability. Figure 3a shows that for the LSTM model, offloading achieves lower average latency as the input utterances are short and their payloads incur only minimal network delay. In the LLM case shown in Figure 3b, which has a higher computational demand than LSTM, the latency reduction from offloading becomes even more pronounced, reaching up to 30 seconds. In both cases, our model predictions closely match the observed latencies.
@@
Key takeaway. Our models generalize to other deep learning workloads by integrating appropriate queuing models.
4.5. Impact of Network Conditions
Figure 4. Latency comparison under varying network bandwidth: (a) RTX4070 as the edge server, (b) A2 as the edge server.
@@ 4.5. Impact of Network Conditions
Figure 4a compares on-device processing on TX2 and Orin Nano with offloading to the RTX4070 GPU under varying network bandwidths. It shows that the performance crossover point depends on both bandwidth and the performance gap between the device and the edge. For the lower-performance TX2, offloading becomes advantageous at 5 Mbps, while the faster Orin Nano requires a higher bandwidth (15 Mbps) for offloading to be beneficial. Figure 4b shows that even for a single device, multiple crossover points may exist depending on bandwidth and the device’s performance. At 5 Mbps, TX2 in its top three frequency levels outperforms offloading, but as its performance drops beyond those levels, offloading becomes advantageous. In all cases, our models closely predict the crossover points, showing their accuracy in capturing the tradeoff between the computation speedup from edge offloading and network delay.
@@
Key takeaway. At high network bandwidths, edge offloading becomes favorable as its computational advantage outweighs the smaller network delays.
4.6. Impact of Collaborative Processing
Figure 5. Latency comparisons of execution strategies under different: (a) collaborative processing configurations, (b) request rates, and (c) numbers of co-located applications.
@@ 4.6. Impact of Collaborative Processing
To validate our models’ ability to capture collaborative (i.e., split) processing behavior, we split the execution of MobilNetV2 across Orin Nano and A2 over a 50 Mbps network. We evaluate three split points (SPs) where the model is progressively offloaded to the edge and compare them to full on-device execution. For each split point, the early neural network layers are executed on the device, and the intermediate output and remaining computation are offloaded to the edge. To compute the end-to-end latency, we use the combined analytic model discussed in Sec 3.5. Figure 5a plots the observed latencies of each configuration alongside model predictions. As shown, the latency increases with the amount of compute offloaded. This is because the intermediate results of later layers grow in size, adding to the network transfer overhead. As seen, for MobileNetV2 and the chosen network configuration, local processing is faster than split processing. Note that other models will exhibit other types of behaviors with collaborative processing, which we omit due to space constraints. Across all split points, the model-predicted latencies closely align with observed values.
@@
Key takeaway. Our models can be extended to accurately predict the performance of device-edge collaborative processing.
4.7. Impact of Request Rate
@@
Figure 5b shows how the request rate $\lambda$ affects offloading decisions for the MobileNetV2 workload under 10 Mbps and 20 Mbps networks, illustrating two distinct behaviors depending on the network bandwidth. At 10 Mbps, Orin Nano outperforms offloading across all RPS. However, at 20 Mbps, a performance crossover occurs: on-device processing is faster at low load (RPS $\leq$ 60), while offloading becomes more efficient beyond that point. This result validates the competing scaling effects of $\lambda$ discussed in Lemma 3.1. Although both the network queuing delays (for offloading) and processing queuing delays grow with $\lambda$, the rate at which the delays increase differs depending on the hardware, workload, and network configurations. At 10 Mbps, the network queuing delay grows faster than the device queuing delay as $\lambda$ increases, favoring on-device processing. Conversely, at 20 Mbps, device queuing delays become dominant at higher $\lambda$, shifting the advantage toward offloading. Overall, the predicted latencies of our models closely match the observed latencies within confidence intervals, demonstrating their accuracy.
@@
Key takeaway. As load increases, on-device processing becomes more advantageous under slower networks, while offloading provides lower average latency on faster networks.
4.8. Impact of Multi-Tenancy